OpenStudy (anonymous):

The Quick Plumbing Supply sells a set of faucets for $79.75. If a 45% markup based on cost was used, what was the cost of the faucets to Quick Plumbing? (Round to the nearest cent.) $

OpenStudy (anonymous):

I would say $50

OpenStudy (anonymous):

x*1.45=79.75

OpenStudy (anonymous):

x=79.75/1.45 x=55

OpenStudy (anonymous):

@princess19 what was your other question?

OpenStudy (chris911):

is it $115.64 ??

OpenStudy (anonymous):

Its 55.00